An aluminium bar 600mm long, with diameter 40mm, has a hole drilled in the center of the bar. The hole is 30mm in diameter and is 100mm long. If the modulus of elasticity for the aluminium is 85GN/m2, calculate the total contraction on the bar due to a compressive load of 180kN.
1
Expert's answer
2020-06-29T08:29:09-0400

Given:

For aluminum bar

compressive Load (P)=180KN

modulus of elasticity (EAL)=85GN/m2

Total Length(L) = 600 mm

Diameter aluminium bar (do) = 40 mm

Hole diameter (di)= 30 mm

Rest bar lenght (L1)=(600-100)=500 mm

Hole length (L2)= 100 mm

Find the total contraction on the bar due to a compressive load .

now, free body Diagram



Now,

"Area\,of\,crosssection\,without\,hole\,(A_{1})=\frac{\pi }{4}\times d_{i}^{2}=\frac{\pi }{4}\times 40^{2}=1256.63\,mm^{^{2}}"


"Area\,of\,crosssection\,with\,hole\,(A_{2})=\frac{\pi }{4}\times (d_{o}^{2}-d_{i}^{2})\\"


"\therefore" "A_{2}=\frac{\pi }{4}\times (40^{2}-30^{2})=549.77\,mm^{^{2}}"


Now,


"Total\,contraction\,=\frac{P\times L_{1}}{A_{1}\times E}+\frac{P\times L_{2}}{A_{2}\times E}=\frac{p}{E}\times (\frac{L_{1}}{A_{1}}+\frac{L_{2}}{A_{2}})"


So,

"Total\,contraction\,=\frac{180\times 10^{3}N}{85\times 10^{3}N/m^{2}}\times \left ( \frac{500}{1256.63} +\frac{100}{549.77}\right )=1.228mm"